OpenStudy (gabby1214):

write an equation in point slope form for the line through the given points with the given slope 1.(8,3); m=6 2.(-10, -6) ; m=-5/8 3.y=3x+7; (2,10) 4. y-3=8/3 (x+2); (-2,3)

pooja195 (pooja195):

\[\huge~\rm~\bf \color{purple}{ y-y_1=m(x-x_1)}\]where m=slope (x1,y1)

pooja195 (pooja195):

Take #1 1.(8,3); m=6 plug it into the equation what do you get?